The ratio girls to boys in a school is 2:3 if 4/7 of the boys are in a drama club , and the remaining 90 boys are not , how many girls are in the school ?
Answer by htmentor(1343)

Girls to boys, G/B = 2/3
4/7B are in drama club. Therefore 3/7B are not in drama club, which is 90 boys.
So (3/7)B = 90 -> B = 630/3 = 210
Thus the number of girls is G = (2/3)*210 = 140